OpenStudy (anonymous):

does this have one solution, no solution, or infinite solutions: 7+2(x-6)=-3+2(x-1)

OpenStudy (lynfran):

7+2(x-6)=-3+2(x-1) 7+2x-12=-3+2x-2 -5+2x=-5+2x -5+5=2x-2x 0=0

OpenStudy (lynfran):

so infinite many solution

OpenStudy (anonymous):

thanks

OpenStudy (lynfran):

welcome